### v1

# SpecOffload: Unlocking Latent GPU Capacity for LLM Inference on Resource-Constrained Devices

Abstract
Efficient LLM inference on resource-constrained devices presents significant challenges in compute and memory utilization. Due to limited GPU memory, existing systems offload model weights to CPU memory, incurring substantial I/O overhead between the CPU and GPU. This leads to two major inefficiencies: (1) GPU cores are underutilized, often remaining idle while waiting for data to be loaded; and (2) GPU memory has low impact on performance, as reducing its capacity has minimal effect on overall throughput. In this paper, we propose SpecOffload, a high-throughput inference engine that embeds speculative decoding into offloading. Our key idea is to unlock latent GPU resources for storing and executing a draft model used for speculative decoding, thus accelerating inference at near-zero additional cost. To support this, we carefully orchestrate the interleaved execution of target and draft models in speculative decoding within the offloading pipeline, and propose a planner to manage tensor placement and select optimal parameters. Compared to the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×. Our code is available at https://github.com/MobiSense/SpecOffload.
1 Introduction
As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation datageneration , form processing formprocessing , and data wrangling datawrangling . These tasks are characterised by LLMs conducting offline inference in batches over large number of tokens. For instance, corporations need to process all archives of financial documentation, whilst individuals want to construct knowledge repositories from accumulated materials. In such workloads, higher inference throughput (the number of tokens generated divided by total generation time, token/s) translates into lower total completion time, hence is the key metric.
Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory constrains. For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) Mixtral8x22b requires at least four NVIDIA H100 (80GB) GPUs. Offloading is one of the mainstream solutions to memory-constrained inference, transferring most model parameters to more economical, capacious CPU memory and reloading them to GPU memory only when computation demands. There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification quantization ; pruning ; sparsification , etc., which are orthogonal and can be applied on top of offloading. Our focus is designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
However, the existing offloading work does not utilize GPU resources well. During offloading, generating each token requires reloading most model parameters from CPU memory to GPU memory for execution deepspeed ; fastinference_offloading ; Klotski . Yet I/O speeds substantially lag behind GPU computational capabilities. For instance, under typical NVIDIA RTX 4090, PCIe 4.0×16 conditions, loading a single FFN layer of the Mixtral 8×22B decoder from CPU to GPU consumes 240ms, while the actual computation on GPU requires merely 0.1ms. Consequently, total inference time is mainly determined by parameter loading time, leaving GPU resources severely underutilized.
Figure 1: GPU core utilization of SOTA methods during decoding phase. Settings: Mixtral 8x7B, Env #1, SummEval dataset, details in § 5.1.
Figure 2: Impact of GPU memory on throughput during decoding phase. Settings: SummEval dataset, details in § 5.1.
To highlight the inefficiencies of existing approaches, we perform a detailed analysis of GPU core and memory utilization. We find that:
  • Underutilization of GPU cores. As shown in Figure 2, during the decoding phase, the average GPU core utilization of existing methods is only 13% at most. This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameters loading. To alleviate this issue, existing methods increase the batch size to amortize I/O overhead—by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski . However, due to the limitation of GPU memory capacity or CPU computational capabilities, scalability of the batch size remains inherently limited. The maximum batch size achieved by the SOTA scheme in Figure 2 is only 64, insufficient to bridge the huge gap between I/O and GPU latency (even with this batch size, the gap remains over 10×).
  • Marginal utility of GPU memory. When models far exceed GPU memory capacity, throughput remains almost unchanged even when memory usage is greatly reduced during decoding phase. We conducted experiments with FlexGen Flexgen , the most effective solution among SOTA approaches. FlexGen offloads attention computations to CPU while computing FFN layers on GPU during the decoding phase. As illustrated in Figure 2, reducing memory usage by over 5.42× for Mixtral 8×7B models led to only a 13% drop in throughput; similarly, a 2.89× reduction for Mixtral 8×22B models caused just a 5% decline. This phenomenon arises because the total volume of data to be loaded remains nearly constant, as the model size vastly exceeds the available GPU memory—leaving little room for variation even when accounting for partical parameters that can reside permanently in memory. For instance, generating a single token necessitates loading all (56) FFN layers for the leftmost blue point in Figure 2, compared to 53 layers for the rightmost blue point. Consequently, memory reduction precipitates only negligible changes in total inference time and throughput.
To harnesses GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD). SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass blockwise ; FastInference ; ACloseLook .
SpecOffload embeds SD into the offloading workflow with nearly zero overhead. The key idea lies in the following two aspects:
  • Computing draft model during GPU core idleness: SD requires draft model to generate multiple candidate tokens in advance. Given the substantial idle periods prevalent in existing offloading frameworks, these intervals can be utilized for completing the draft model’s computational tasks.
  • Storing draft model use "low-yield" GPU memroy: SD requires loading a draft model into memory for draft generation. We can repurpose "low-yield" memory allocations from existing frameworks to store draft model parameters and its caches instead. For instance, as shown in Figure 2, extracting 17GB of "low-yield" memory allocation provides sufficient capacity for a draft model such as Mistral 7B mistral7b to operate normally within the GPU at a small batch.
To support this, SpecOffload designs a comprehensive framework to better utilize both the computational and memory resources of the GPU. SpecOffload determines tensor distribution between GPU and CPU memory through offline Adaptive Tensor Placement (§ 4.2), dynamically schedules computational tasks via the online ParaSpec Planner (§ 4.3), and implements parallel pipelined execution of I/O and computation using the Interleaved Batch Pipeline (§ 4.1).
Our contributions are as follows:
  • We conduct a quantitative analysis of GPU resource utilization in representative scenarios and identify key limitations in SOTA frameworks——underutilization of GPU cores and marginal utility of GPU memory, thus reveal a novel perspective for enhancing offloading performance.
  • By designing sub-layer model decomposition and fine-grained scheduling compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
  • To evaluate SpecOffload, we benchmark it against HuggingFace Accelerate accelerate , DeepSpeed-FastGen deepspeed , FlexGen Flexgen , and Fiddler fiddler . Results demonstrate that our system achieves throughput improvements of 4.69×, 4.71×, 2.54×, and 4.04× over these baselines, respectively.
2 Background and Related Work
2.1 Generative LLM Inference
Large language models consist of stacked Transformer layers attentionisallyouneed . During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated token and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
2.2 Speculative Decoding
Speculative decoding (SD) is a method for accelerating LLM inference. It adheres to a “Draft-then-Verify” framework: at each decoding step, a lightweight draft model initially proposes multiple candidate tokens (e.g., $(\hat{w}_{1},\hat{w}_{2},\hat{w}_{3},\hat{w}_{4}$)), which are collectively verified by a larger target model in a single forward pass. Only the valid subset $({w}_{1},{w}_{2})$ is accepted, after which the target model resumes decoding by independently generating the subsequent token $w_{3}$ blockwise ; FastInference ; ACloseLook . This approach enables the target model to generate multiple tokens per inference step. To further enhance the efficiency of SD, prior research has predominantly explored two avenues: the design of more effective draft models distillspec ; draft-verify ; onlineSD and draft structures medusa ; specinfer ; specexec . However, traditional single-batch SD is not well-suited for integration with offloading, as the computations of the draft and target models must be executed sequentially. As a result, the GPU resources still remain underutilized during the target model’s verification phase. By introducing a dual-batch rotation strategy, SpecOffload enable the verifying and drafting to run concurrently, allowing SD to be seamlessly embedded into the offloading pipeline, while better utilizing GPU resources.
2.3 Offloading in LLM Inference
Offloading is one of the predominant solutions for enabling LLM inference under GPU memory constraints. It entails relocating certain model parameters from the expensive, limited GPU memory to the more cost-effective and abundant CPU memory deepspeed ; fiddler ; powerinfer ; moe-infinity ; fastdecode ; fastinference_offloading . I/O constraints create the primary bottleneck in offloading, as data transfer latency between CPU memory and GPU memory far exceeds the GPU’s computation time. Consequently, GPU resources remain underutilized.
In throughput-oriented scenarios, existing methods typically increase the batch size to amortize I/O costs. A prevalent strategy involves altering the model’s execution pattern from row-wise to column-wise, allowing each layer’s parameters to be loaded once and reused across multiple batches, thereby reducing the per-layer I/O burden Flexgen ; Klotski . However, this approach is constrained by the available GPU memory and the overhead I/O of KV cache. More recent studies demonstrate that offloading attention computation to the CPU can eliminate KV cache I/O in decoding phase moe-lightning ; neo ; hetegen ; PACT'24 , but in doing so, CPU compute limitations cap batch scalability. Thus, GPU resources remain significantly underutilized. In this work, by interleaving the draft model’s workload into the GPU’s idle periods between the target model’s layer-wise computations, we improve GPU core utilization by 4.49×.
3 System Overview
Figure 3: System overview of SpecOffload.
In this work, we propose SpecOffload. As shown in Figure 3, it employs a two-phase architecture where offline tensor placement and online scheduling collectively determine a unified pipeline across GPU and CPU.
During the offline phase, the target and draft models are deployed across heterogeneous memory hierarchy. SpecOffload automatically evaluates the hardware ecosystem, measuring CPU and GPU memory capacities, computational performance of CPU and GPU cores, and the bandwidth of data transfer channels between them. These hardware specifications, along with the configurations of models (➀), are input to the Adaptive Tensor Placement (§ 4.2) to determine the initial parameter allocation strategy (➁).
During the online phase, the hardware configuration and batched inputs (➂) are provided to the ParaSpec Planner (§ 4.3), which, based on input length and characteristics, computes a fine-grained pipeline execution plan, including the batch sizes for target model in prefill and decoding phase, batch size for draft model, and the number of candidate tokens to generate (➃).
The scheduling results from both offline and online phases (➄,➅) collectively determine the Interleaved Batch Pipeline (§ 4.1). The pipeline consists of three main threads: GPU computation, CPU computation, and GPU-CPU I/O. Parameter residency within GPU and CPU memory dynamically adapts in response to the progression of the I/O thread (➆).
4 Method
4.1 Interleaved Batch Pipeline
Motivated by the differing computational demands of the prefill and decoding phases in LLM inference, we introduce the Interleaved Batch Pipeline—a phase-specific pipeline design. During the prefill stage, the target model computation dominates the GPU runtime and we perform additional memory management before completion. In the decoding stage, we embed speculative decoding into the pipeline by finely interleaving the computations of the two models to fully utilize the GPU cores.
4.1.1 Prefill Phase
While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen Flexgen , we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding. In order to minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
4.1.2 Decoding Phase
During the decoding phase, we build upon the original offloading framework by repurposing the low-yield GPU memory to store the draft model and leveraging GPU idleness to execute it. To enable this, we redesign the entire pipeline at both the model and computation levels. The decoding phase pipeline is illustrated in Figure 4. In summary, our model-level design transforms conventional single-batch speculative decoding into a dual-batch interleaved scheme to facilitate parallel execution. At computation-level, the draft model’s workload is finely interleaved into the GPU idle periods between the layer-wise computations of the target model.
Figure 4: Schematic of the decoding pipeline. At model-level, while the target model validates Batch 1, the draft model concurrently generates tokens for Batch 0 (in time slot $t_{n}$); the two models then alternate batches (in time slot $t_{n+1}$). At computation-level, the target model’s parameters are computed layer by layer. GPU, IO, and CPU are orchestrated to perform distinct, parallelized tasks.
From model-level, the decoding phase involves two batches being processed alternately by the target and draft model. Conventional speculative decoding adopts a single-batch Draft-then-Verify paradigm, as the computations of the draft and target models must be executed sequentially. Consequently, the GPU resources left idle during the target model’s verification stage still remain unused, while the system suffers from additional overhead caused by switching between models.
To overcome these limitations, we propose a dual-batch interleaved design that enables true model-level parallelism. As shown on the left side of Figure 4, in time slot $t_{n}$, while the target model verifies Batch 1, the draft model concurrently generates speculative tokens for Batch 0 ($\hat{w}_{1}$, $\hat{w}_{2}$, $\hat{w}_{3}$, $\hat{w}_{4}$). Once both tasks are completed, the roles switch. In time slot $t_{n+1}$, the target model validates Batch 0 (${w}_{1}$, ${w}_{2}$, ${w}_{3}$, ${w}_{4}$), and the draft model proceeds with Batch 1. This alternating batch rotation continues until the generation is complete. This batch interleaving primarily enables parallel execution of drafting and verification.
From computation-level, the decoding pipeline involves coordination among three threads: GPU computation, CPU computation, and I/O. The right side of Figure 4 provides an illustrative example. Each layer of the target model is fine-grainedly partitioned into attention and FFN components. For clarity, lightweight components such as normalization layers are omitted from the figure, as their parameter size and computational cost are negligible. As illustrated, for the majority of the inference time, the CPU performs attention computation (current batch), the system transfers same layer’s FFN parameters from CPU memory to GPU memory (current batch), and the GPU executes draft model computation (another batch), all in parallel. Upon completion of attention computation on the CPU, the intermediate activations are transferred to the GPU. Finally, once both the parameters and activations are available on the GPU, the remaining computations are quickly completed. After completing all layers of the target model (the end of a round of parallel drafting and verification at model level), the target and draft models exchange their current batches.
Details such as where model parameters are computed, how I/O is handled, and batch configurations are determined in the § 4.2 and § 4.3.
4.2 Adaptive Tensor Placement
SpecOffload introduces a novel design for heterogeneous models (target and draft models), by jointly managing the parameter placement across GPU memory, CPU memory, and disk tiers, thus enabling efficient speculative decoding. Adaptive Tensor Placement strategy intelligently assigns tensors to different memory tiers based on real-time resource availability and the current computational task, optimizing memory utilization and mitigating I/O bottlenecks.
We establish tensor prioritization hierarchically by sub-layer, categorizing based on both functional type (attention, KV cache, FFN) and computational phase. Tensors required by the current and next layers of the target model are assigned the highest priority and are preferentially placed in GPU memory. Draft model and its cache are also treated as high-priority and retained in GPU memory during decoding phase. If GPU memory capacity permits, additional parameters are pinned to further reduce I/O overhead. Remaining tensors are offloaded to CPU memory with moderate priority, leveraging its high bandwidth and low latency, as well as its ability to support certain computations. If CPU memory is exhausted, parameters are further offloaded to disk. When CPU memory is sufficient, $pin\_memory()$ is employed to accelerate GPU-CPU data transfer. A dynamic memory management mechanism is employed to avoid cross-tier memory swaps, ensuring that only CPU memory interfaces with both GPU memory and disk.
The core of the dynamic memory management mechanism is prefetching, which overlaps I/O with computation. For example, while computing attention of layer $i$, GPU memory preloads FFN of the same layer from CPU memory, and concurrently, CPU memory prefetches the parameters of layer $i+1$ from disk. Dedicated placeholders are reserved in GPU & CPU memory for prefetched tensors.
4.3 ParaSpec Planner
Interleaved Batch Pipeline section (§ 4.1) outlines our pipeline strategy; however, key parameters—such as the batch sizes of target model during prefill and decoding phase, batch size of draft model, generated draft token number, require careful tuning. To address this, we propose ParaSpec Planner, a parameter specialization module that selects optimal configurations for a given input.
Planning Goal. ParaSpec Planner aims to maximize model inference throughput on a given hardware configuration. Throughput is determined by two factors: the total number of tokens generated per batch inference, denoted as $\tilde{N}_{generated}$, and the corresponding generation latency, $T_{generation}$. On consumer-grade hardware, the primary system constraints lie in GPU memory capacity. Therefore, we formulate the problem as a constrained optimization task as follows:
$\max~{}throughput$ $=\max~{}\frac{\tilde{N}_{generated}}{T_{generation}}$
$s.t.~{}~{}gpu~{}peak~{}memory~{}$ $\leq~{}gpu~{}mem~{}capacity$ (1)
Generated Tokens. The total number of generated tokens, $\tilde{N}_{{generated}}$, is the sum of tokens $\tilde{n}_{{generated}}$ produced over $n_{iter}$ iterations for a batch of size $bs$. However, in our system, speculative decoding introduces randomness, causing the number of tokens generated per input in each iteration to become a random variable. We use the expected value to represent the average number of tokens that pass verification in each iteration.
$\tilde{N}_{generated}=\sum_{bs}\sum_{n_{iter}}\tilde{n}_{generated}=bs\times n_{iter}\times\mathbb{E}[n_{generated}]$ (2)
Inference Latency. The inference latency $T_{generation}$ is determined by the degree of parallelism in the inference pipeline. Due to architectural differences between the prefill and decoding phases in SpecOffload, their latencies must be modeled separately, $T_{generation}=T_{prefill}+T_{decoding}$. Since computation in the prefill phase is primarily GPU-bound, its execution time is independent of batch size and instead depends on the number of computation steps required.
$T_{prefill}=\left\lceil\frac{bs}{bs_{prefill}}\right\rceil\times T_{target,prefill}^{GPU}$ (3)
During the decoding phase, SpecOffload performs two primary tasks: draft generation for one batch and verification for another. The overall latency is determined by the longer of the two.
$T_{decoding}=\max(T_{target,decoding},~{}~{}T_{draft})$ (4)
Memory Constraints. GPU memory constraints can likewise be decomposed into those for the prefill and decoding phase. In each phase, the combined memory footprint of model parameters, intermediate activations, and KV cache must not exceed the available GPU memory. In the prefill phase, GPU memory consumption is primarily composed of two parts: the parameter size of the target model, and the KV cache required.
$V_{prefill}=V_{target,prefill}+V_{target,KVcache}$ (5)
Similarly, in decoding, GPU memory usage consists of the main model parameters, the draft model parameters, and the KV cache used by the draft model.
$V_{decoding}=V_{target,FFN}+V_{draft}+V_{draft,KVcache}$ (6)
More detailed derivations, please refer to the Appendix A.1. Before using the ParaSpec Planner, a profiling program must be run on the target hardware to collect performance characteristics. However, due to the challenges of hardware measurement, OS-induced variability, and the uncertainty in draft token validity introduced by speculative decoding, such measurements may not fully reflect the actual behavior of SpecOffload during execution. Consequently, while ParaSpec Planner can produce high-quality parameter configurations, further fine-tuning may still be required to achieve optimal performance.
5 Evaluation
Table 1: Hardware Configurations.
Env #1 Env #2
GPU RTX 4090 RTX 4090
VRAM 24G 24G
PCIe Gen3 x 16 Gen4 x 16
CPU i9-10980XE EPYC 7542
DRAM 256G 448G
Table 2: Dataset Configurations.
HumanEval humaneval C-Eval ceval SummEval summeval SAMSum samsum
$S_{avg}$ 157.54 165.46 503.02 168.10
$S_{max}$ 437 483 783 1144
$S_{std}$ 72.46 103.18 138.68 120.53
Task Coding Exam Summarization
5.1 Experimental Setup
Implementation. We implement SpecOffload on top of Hugging Face Transformers v4.47.1transformers . We implement pipeline using multiprocessing with shared memory for inter-process vector communication. More details is provided in Appendix A.2.
Models. We evaluate SpecOffload using two popular and open-source models: Mixtral-8x7B Mixtral8x7b and Mixtral-8x22B Mixtral8x22b . They have 46.7B and 141B parameters in bfloat16 precision. The draft model for speculative decoding is Mistral-7B mistral7b . Although not evaluated, SpecOffload can support other models compatible with Transformers’ model classes.
Hardware. We evaluate SpecOffload in two different evironments, as shown in Table 2.
Datasets. We use most common LLM benchmarks with different prompt length distributions and tasks, as shown in Table 2. HumanEval dataset humaneval released by OpenAI includes 164 programming problems; C-Eval dataset ceval is a comprehensive Chinese evluation suite includes 13948 questions; SummEval dataset summeval inlcudes 100 news article from the CNN/DailyMail; SAMSum dataset samsum contains about 16k messenger-like conversations with summaries.
Baselines. We compare against 4 baseline systems, all designed to address GPU memory limitations.
  • Hugging Face Accelerate accelerate supports offloading weights of some layers based on the device map. We use its version 1.5.2. Hereinafter referred to as Accelerate.
  • DeepSpeed Zero-Inference deepspeed supports offloading the whole weights to the CPU or disk. We use its version 0.16.1. Hereinafter referred to as DeepSpeed.
  • FlexGen Flexgen employs a "zig-zag" inference schedule to increase throughput.
  • Fiddler fiddler strategically utilizes both CPU and GPU resources for MoE model inference.
Additionally, except for FlexGen, all other baselines natively support Mixtral. We adapted FlexGen to Mixtral while adhering to its original offloading strategy.
Metrics. Throughput (token/s) is calculated as the number of tokens generated divided by total genration time (prefill time + decoding time).
5.2 End-to-end Throughput
Figure 5: End-to-end comparison between SpecOffload and baselines in different scenarios.
Figure 5 illustrates the end-to-end throughput of five approaches across two environments and three datasets. In Env #1, using the Mixtral 8×7B model, SpecOffload attains an average speedup of 2.53× over the strongest baseline, FlexGen; in Env #2, with the Mixtral 8×22B model, it achieves an average speedup of 2.54×. The superior throughput of SpecOffload is attributed to the embedding of speculative decoding into the offloading pipeline. Empirical results demonstrate that computing and storing the draft model on the GPU memory confers greater performance gains than offloading either the target model’s parameters or its cache.
The performance of our method on additional datasets is provided in the Appendix A.3.1. More Details on the policy setups and effective batch sizes can be found in Appendix A.3.2.
5.3 Effectiveness Analysis
Table 3: Detail Runtime breakdown (seconds)."P" and "D" denote Prefill and Decoding, respectively. Compute(G,T) and Compute(G,D) denote the GPU computation time for the target and draft models, respectively, while Compute(C) represents the target model’s computation time on the CPU. Cache(G→C) indicates the time to transfer the KV cache from GPU memory to CPU memory.
Phase Total Compute(G,T) Compute(G,D) Compute(C) Weight(R) Cache(G→C)
8×7B, Env #1 P 183.28 79.62 0 0 123.48 39.05
D 569.21 35.34 489.02 531.23 236.2 0
8×22B, Env #2 P 280.42 42.22 0 0 166.45 91.06
D 794.26 27.34 345.93 746.38 262.64 0
Figure 6: Decoding phase GPU core utilization.
Figure 7: Decoding phase memory consumption.
We employ NVIDIA Nsight nsight to monitor GPU core utilization and memory consumption during the decoding phase of Mixtral 8×7B in Env #1 on the SummEval dataset. As depicted in Figure 7, the average core utilization reaches 58.67%, attains 8.14×, 7.15×, 4.49×, and 8.24× higher than Accelerate, DeepSpeed, FlexGen, and Fiddler, respectively. The elevated core utilization can be attributed to the residency of the draft model in GPU memory, enabling continuous computation.
Figure 7 reveals a periodic pattern in the draft model’s GPU memory consumption. Each cycle lasts approximately 28 seconds, characterized by a gradual increase in memory usage followed by a sharp drop and a 2-second idle window. This aligns with the behavior observed in Figure 7, where the draft model performs computation for 26 seconds and remains idle for 2 seconds awaiting the next batch. More detailed GPU memory use in Appendix A.3.3.
We show the runtime breakdown of Mixtral 8x7B in Env #1 and 8x22B in Env #2 in Table 3 on SummEval dataset. We disable overlapping and profile the time used for GPU, I/O and CPU. Results show that our method effectively overlaps compute and I/O.
5.4 Ablation Study
Table 4: Ablation study of proposed techniques on SummEval dataset. The gray tuple denotes a policy (prefill batch size, decoding batch size, draft batch size, draft max new tokens).
All optimizations No policy search Serial SD No SD
8x7B 24.743 (80, 192, 8, 8) 15.624 (50, 256, 5, 2) 17.048 (80, 192, 40, 8) 12.369 (80, 256, x, x)
8x22B 5.911 (16, 64, 8, 8) 3.486 (16, 32, 6, 6) 4.146 (16, 64, 32, 8) 1.698 (16, 80, x, x)
We then isolate the improvement brought by each individual technique. Table 4 lists the throughput SpecOffload can achieve if disabling one technique at a time. On Mixtral 8x7B, we choose target model prefill batch 80, decoding batch size of 192. And due to the design of rotating batches, the total batch size is 192 $\times$ 2 = 384. Draft model batch of 8 and generates 8 tokens per iteration. "No policy search" illustrates the performance of a random strategy, showing the importance of a good policy. We also decoupled speculative decoding from the tightly integrated pipeline into a loosely coupled serial mode. Results show that embedding speculative decoding into the pipeline is beneficial, as the naive combination of serial speculative decoding and offloading incurs additional I/O overhead from the draft model and its KV cache. Removing speculative decoding highlights its substantial performance gains. Additional ablation studies are provided in Appendix A.3.4.
5.5 Load to disk
Figure 8: Throughput comparison of Mixtral 8×22B with and without disk offloading. No Disk corresponds to execution in Env #2, which has sufficient CPU memory, while Disk corresponds to execution in Env #1, where GPU memory is limited.
We further conducted experiments in Env #1 by extending the offloading to disk. The disk read and write speeds are 3.5GB/s and 1.7GB/s, respectively. Env #1 offers a memory capacity of 256GB, which is insufficient to accommodate the complete Mixtral 8×22B model (141B parameters). As shown in Figure 8, under these memory-limited conditions, load to disk SpecOffload attains 29.3% of its original throughput.
6 Conclusion
We identify two key inefficiencies in existing offloading frameworks for LLM inference: underutilization of GPU cores and marginal utility of GPU memory. To address these, we propose SpecOffload, which embeds speculative decoding into offloading with virtually zero overhead by leveraging idle GPU time and "low-yield" GPU memory. Experiments show up to 2.54× throughput gains over the best baseline, demonstrating the effectiveness of our approach for high-throughput LLM inference on resource-constrained devices.
Appendix A Technical Appendices and Supplementary Material
A.1 ParaSpec Planner
Planning Goal.
ParaSpec Planner aims to maximize model inference throughput on a given hardware configuration. Throughput is determined by two factors: the total number of tokens generated per batch inference, denoted as $\tilde{N}_{generated}$, and the corresponding generation latency, $T_{generation}$. On consumer-grade hardware, the primary system constraints lie in GPU memory capacity. Therefore, we formulate the problem as a constrained optimization task as follows:
$\max~{}throughput$ $=\max\frac{\tilde{N}_{generated}}{T_{generation}}$
$s.t.~{}~{}gpu~{}peak~{}memory~{}$ $\leq~{}gpu~{}mem~{}capacity$ (7)
Generated Tokens.
The total number of generated tokens, $\tilde{N}_{{generated}}$, is the sum of tokens $\tilde{n}_{{generated}}$ produced over $n_{{iter}}$ iterations for a batch of size $bs$ In conventional decoding without speculation, each input generates exactly one token per iteration, making $\tilde{n}_{{generated}}$ a constant. However, in our system, speculative decoding introduces randomness, causing the number of tokens generated per input in each iteration to become a random variable. As a result, $\tilde{n}_{{generated}}$ cannot be expressed deterministically and is instead characterized by its expected value. Let $\tilde{n}_{{generated}}$ denote the actual number of tokens generated for a single input, then $\tilde{n}_{{generated}}=\mathbb{E}[n_{{generated}}]$, as shown in Equation 9.
$\tilde{N}_{generated}$ $=\sum_{bs}\sum_{n_{iter}}\tilde{n}_{generated}$ (8)
$=bs\times n_{iter}\times\mathbb{E}[n_{generated}]$ (9)
To characterize the distribution of $\tilde{n}_{\text{generated}}$, we model the speculative decoding process. In each iteration, the draft model generates a candidate sequence of $n_{\text{cand}}$ tokens, which is then verified by the target model. The target model returns the longest correct prefix of the candidate sequence and subsequently generates one additional correct token. The number of tokens correctly predicted by the draft model ranges from 00 to $n_{\text{cand}}$, so $\tilde{n}_{\text{generated}}$ follows a distribution over the set $\{1,\dots,n_{\text{cand}}+1\}$.
We assume that the probability of the draft model correctly predicting a single token is $p$, and that these predictions are independent across positions. Under this assumption, the probability that the main model accepts exactly $k$ tokens is given by the probability that the first $k-1$ tokens are correct and the $k$th is incorrect, as shown in Equation 10. If $k=n_{\text{cand}}+1$, it corresponds to the draft model correctly predicting the entire candidate sequence, this probability distribution is formalized in in Equation 11.
$\mathbb{P}[n_{generated}=k]=p^{k-1}\cdot(1-p_{cand}),\quad k=1,\cdots,n_{cand}$ (10)
$\mathbb{P}[n_{generated}=k]=p^{k-1},\quad k=n_{cand}+1$ (11)
The expected value $\mathbb{E}[n_{\text{generated}}]$ is derived in Equation 12. Thus, the total number of tokens generated by the model, $\tilde{N}_{\text{generated}}$, is expressed as a function of $bs$, $n_{\text{iter}}$, $n_{\text{cand}}$, and $p$.
$\mathbb{E}[n_{generated}]$ $=\sum_{k=1}^{n_{cand}+1}k\cdot\mathbb{P}[n_{generated}=k]$
$=\frac{1}{1-p}[n_{cand}p^{n_{cand}+2}-(n_{cand}+1)p^{n_{cand}+1}+1]$ (12)
Inference Latency.
The inference latency $T_{\text{generation}}$ is determined by the degree of parallelism in the inference pipeline. As SpecOffload exhibits distinct behaviors in the Prefill and decoding stages, their latencies must be computed separately, in Equation 13.
$T_{generation}=T_{prefill}+T_{decoding}$ (13)
In the Prefill stage, loading the full KV cache for all $bs$ inputs would exceed GPU memory capacity. Therefore, SpecOffload partitions the batch into small Prefill batch $bs_{{prefill}}$ Since computation in the Prefill stage is primarily GPU-bound, its latency is independent of the Prefill batch size and instead determined by the number of iterations required, as formalized in Equation 14.
$T_{prefill}=\left\lceil\frac{bs}{bs_{prefill}}\right\rceil\times T_{target,prefill}^{GPU}$ (14)
In each iteration, the processing time per $bs_{\text{Prefill}}$ is primarily determined by parameter I/O ($T^{C2G}_{para}$) and computation ($T^{GPU}_{target,comp}$), with I/O time significantly exceeding computation time in the offloading scenario, as shown in Equation 15.
$T_{target,prefill}^{GPU}=T^{C2G}_{para}+T^{GPU}_{target,comp}\approx T^{C2G}_{para}$ (15)
In the decoding stage, SpecOffload performs two primary tasks in parallel: draft generation for one batch and verification for another. The overall latency is thus determined by the slower of the two tasks in Equation 16.
$T_{decoding}=\max(T_{target,decoding},~{}~{}T_{draft})$ (16)
The draft generation task incurs a latency equal to the time required to execute the draft model inference entirely on the GPU. Similarly, due to memory constraints, the draft model must also divide each batch into smaller sub-batches $bs_{draft}$ for generation, $T_{draft}^{GPU}$ is the time for one-batch generation. Each generation step can be further decomposed into Prefill and decoding stages, as shown in Equation 17.
$T_{draft}$ $=\left\lceil\frac{bs}{bs_{draft}}\right\rceil\times T_{draft}^{GPU}$
$=\left\lceil\frac{bs}{bs_{draft}}\right\rceil\times[T_{draft,prefill}^{GPU}+(n_{cand}-1)T_{draft,decoding}^{GPU}]$ (17)
For the verification task, based on SpecOffload ’s pipeline design, each decoder layer’s FFN computation depends on both the output of the Attention module and the loading of FFN parameters. Attention computation and FFN loading are executed in parallel threads, with the completion time determined by their maximum. Subsequently, the GPU performs FFN computation, which is significantly faster than parameter loading. Therefore, its latency can be expressed as in Equation 18.
$T_{target,decoding}$ $=n_{layer}\times[\max(T_{target,Attention}^{CPU},T_{target,FFN}^{C2G})+T_{target,FFN}^{GPU}]$
$\approx n_{layer}\times[\max(T_{target,Attention}^{CPU},T_{target,FFN}^{C2G})]$ (18)
Importantly, since the Attention module is offloaded to the CPU, its runtime becomes dependent on the sub-batch size, and is modeled accordingly in Equation 19.
$T_{target,Attention}^{CPU}=n_{cand}\times bs\times t_{target,Attention}^{CPU}$ (19)
Memory constraint.
The GPU memory capacity constraint can similarly be decomposed into separate constraints for the prefill and decode phases. In both phases, the total GPU memory consumption—including model parameters, intermediate activations, and KV cache—must not exceed the available GPU memory, which directly impacts the feasible batch size per inference.
In the prefill phase, GPU memory usage primarily consists of two components: the memory footprint of the main model parameters ($V_{\text{main}}$) and the KV cache required during prefill ($V_{m,\text{KVcache}}$). Since the KV cache for all $bs$ inputs would exceed GPU memory capacity, the batch is partitioned into sub-batches of size $bs_{\text{prefill}}$. As a result, the KV cache footprint in the prefill phase only accounts for $bs_{\text{prefill}}$ inputs, as formalized in Equation 20.
$V_{prefill}$ $=V_{target,prefill}+V_{target,KVcache}$
$=V_{target,prefill}+bs_{prefill}\times l_{input}\times v_{target,KVcache}$ (20)
Similarly, GPU memory usage in the decode phase consists of three components: the main model parameters loaded into GPU memory, the draft model parameters, and the KV cache used by the draft model. According to SpecOffload’s offloading strategy, only the MoE FFN parameters from the main model are loaded into GPU memory, whereas the draft model is fully resident in GPU memory. Therefore, the total GPU memory footprint during decoding is characterized in Equation 21. To satisfy the memory constraint, the batch of $bs$ inputs is partitioned into sub-batches of size $bs_{\text{assist}}$, as defined in Equation 22.
$V_{decoding}$ $=V_{target,FFN}+V_{draft}+V_{draft,KVcache}$ (21)
$=V_{target,FFN}+V_{draft}+bs_{draft}\times(l_{input}+n_{generated})\times V_{draft,KVcache}$ (22)
A.2 Implementation
Figure 9: Implementation from the perspective of Interleaved batches.
Our implementation is based on modifications to HuggingFace Transformers [39], version 4.47.1.
To implement the SpecOffload pipeline, we adopt a hybrid parallelism strategy that combines process-level and thread-level parallelism. As shown in Figure 9, the input sequence is split into two interleaved batches, which alternate between draft generation and large-model verification. Each batch is processed on a separate thread, with synchronization managed via inter-thread events. After completing its generation and verification task in each iteration, a thread signals its completion and waits for the other thread to do the same. The next iteration begins only after both threads have finished the current one. This design enables parallel execution of Batch 0’s draft generation and Batch 1’s verification. However, from the perspective of a single batch (Batch 0 or Batch 1), the draft and verify stages remain sequential.
In this design, the draft model performs full-sequence autoregressive inference on GPU during the generation stage, while the large model remains computing on CPU to avoid resource contention. Given that the draft model executes strictly in a sequential manner—processing only one batch at any time—parallel instantiation is unnecessary. To further minimize GPU memory consumption, only a single copy of the draft model is loaded and isolated within a dedicated auxiliary process. Communication between this draft model process and the main process hosting the large model is established via shared memory, enabling low-latency data transfer, while inter-process events are employed to enforce strict execution ordering and synchronization.
Figure 10: Inter-process communication diagram. Orange blocks represent the send/receive modules within each process, while green blocks indicate the inter-process communication modules. "Input" and "Output" are defined with respect to the Draft process.
As shown in Figure 10, inter-process communication is centered around two shared memory regions, with tokens as the primary data exchanged. The Draft process writes generated draft tokens to the output shared memory for consumption by the Main process, and reads verified tokens from the input shared memory region. When a thread in the Main process reaches the verify stage and requires draft tokens, it waits for the draft model to signal availability, then read from shared memory. Upon finishing computation, the draft model signals completion and readiness for the next task using event flags, ensuring proper synchronization and preventing data races or overlap between batches. This tightly coordinated mechanism enables efficient and orderly pipelined execution across model components, while keeping both memory footprint and runtime overhead to a minimum.
A.3 Additional Experimental Results
A.3.1 End-to-end Throughput
Figure 11: End-to-end comparison between SpecOffload and baselines on SAMSum.
In addition to the datasets discussed in the main text, we further evaluated SpecOffload on the SAMSum dataset [38], which has an average input length of 168.1 and a maximum length of 1144. The SAMSum dataset contains about 16k messenger-like conversations with summaries. The results on SAMSum are consistent with those observed on the other datasets. As shown in Figure 11, SpecOffload achieves a throughput of 2.575× that of the best baseline on the Mixtral 8×7B model, and 2.25× on the Mixtral 8×22B model.
A.3.2 Impact of Policy
We present detailed end-to-end throughput data, as shown in Table 5, Table 6, Table 7, Table 8, Table 9, and Table 10, to simulate different scenarios and analyze the impacts of policy on throughput by generating 16 tokens. Due to the large number of GPU hours required to complete all (prefill batch size, decoding batch size, draft batch size, draft max new tokens) combinations using Mixtral 8×22B in Env #2, we evaluated only a subset of possible configurations.
The prefill batch size is a tunable parameter for which the optimal value can be explicitly determined by the scheduling algorithm, as the Prefill stage does not involve speculative decoding and thus is free from probabilistic uncertainty. Under the experimental setup of Table 7, the optimal value is 80. For example, comparing entries 5 and 27 in Table 7—where all other parameters are held constant—the higher throughput of entry 27 is attributed to its more optimal Prefill batch size.
The decoding batch size and draft max new tokens jointly affect the verification latency of the target model. Since the target model’s computation is offloaded to the CPU, the speculative decoding verification cannot achieve the same level of tight serialization as on the GPU. As a result, increasing the batch size or the number of new tokens leads to longer CPU computation time. As illustrated by entries 26–30 and 10, 20, 30, 40 in Table 7, neither a larger batch size nor a higher max new token value consistently yields better performance.
The decoding batch size, draft batch size, and draft max new tokens jointly impact the generation latency of the draft model. Due to GPU memory constraints, the draft batch size is typically limited to a small value. However, since all draft model computations are executed on the GPU, they are highly efficient. This allows the full decoding batch to be processed through a fine-grained, multi-round strategy. As long as the draft model’s token generation time remains below the I/O-bound latency, it does not constitute a performance bottleneck.
The results in Table 10 exhibit a similar pattern. These four parameters are tightly coupled and collectively determine the overall system throughput. Given that our design introduces at least four tunable parameters, finding optimal settings through enumeration or heuristics alone is highly unlikely. This highlights the critical role of the Paraspec Planner in the overall system.
Table 5: Impact of policy on Mixtral 8x7B in Env #1, HumanEval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 80 200 10 8 32.821
2 80 160 6 1 15.869
3 80 160 6 2 20.964
4 80 160 6 4 28.914
5 80 160 6 6 33.711
6 80 160 6 8 33.690
7 80 160 8 1 15.834
8 80 160 8 2 20.940
9 80 160 8 4 29.267
10 80 160 8 6 32.520
11 80 160 8 8 32.776
12 80 160 10 1 15.835
13 80 160 10 2 21.120
14 80 160 10 4 29.499
15 80 160 10 6 32.226
16 80 160 10 8 32.540
17 80 200 6 1 18.736
18 80 200 6 2 24.737
19 80 200 6 4 29.091
20 80 200 6 6 31.641
21 80 200 6 8 33.014
22 80 200 8 1 18.828
23 80 200 8 2 24.813
24 80 200 8 4 30.452
25 80 200 8 6 32.649
26 80 200 8 8 31.884
27 80 200 10 1 18.865
28 80 200 10 2 24.675
29 80 200 10 4 30.363
30 80 200 10 6 32.716
31 80 200 10 8 33.072
32 80 256 6 1 21.166
33 80 256 6 2 26.052
34 80 256 6 4 30.279
35 80 256 6 6 32.325
36 80 256 6 8 32.812
37 80 256 8 1 20.683
38 80 256 8 2 27.123
39 80 256 8 4 31.829
40 80 256 8 6 33.622
41 80 256 8 8 33.247
42 80 256 10 1 20.546
43 80 256 10 2 26.987
44 80 256 10 4 30.679
45 80 256 10 6 34.665
46 80 256 10 8 33.445
Table 6: Impact of policy on Mixtral 8x7B in Env #1, C-Eval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 96 256 8 4 26.489
2 96 288 8 4 28.449
3 96 300 8 4 28.209
4 96 256 6 2 25.363
5 96 256 6 4 27.823
6 96 256 6 6 28.712
7 96 256 6 8 28.531
8 96 256 8 2 25.347
9 96 256 8 4 27.449
10 96 288 6 2 25.254
11 96 288 6 4 28.685
12 96 288 6 6 29.199
13 96 288 6 8 29.385
14 96 288 8 2 26.126
15 96 288 8 4 28.679
16 96 288 8 6 29.329
17 96 300 6 2 24.821
18 96 300 6 4 28.240
19 96 300 6 6 29.134
20 96 300 6 8 30.781
21 96 300 8 2 26.268
22 96 300 8 4 30.652
23 96 300 8 6 31.968
Table 7: Impact of policy on Mixtral 8x7B in Env #1, SummEval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 50 128 5 5 18.937
2 50 128 5 3 19.735
3 50 256 5 5 19.890
4 50 256 5 3 17.560
5 50 256 5 2 15.624
6 80 128 5 1 11.682
7 80 128 5 2 14.509
8 80 128 5 4 19.464
9 80 128 5 6 21.166
10 80 128 5 8 21.531
11 80 128 8 1 11.629
12 80 128 8 2 14.408
13 80 128 8 4 18.321
14 80 128 8 6 16.989
15 80 128 8 8 21.958
16 80 192 5 1 14.764
17 80 192 5 2 16.830
18 80 192 5 4 21.072
19 80 192 5 6 22.029
20 80 192 5 8 22.712
21 80 192 8 1 14.305
22 80 192 8 2 16.757
23 80 192 8 4 21.435
24 80 192 8 6 23.653
25 80 192 8 8 24.732
26 80 256 5 1 14.809
27 80 256 5 2 16.781
28 80 256 5 4 20.441
29 80 256 5 6 21.841
30 80 256 5 8 21.741
31 80 256 8 1 13.822
32 80 256 8 2 16.265
33 80 256 8 4 17.243
34 80 256 8 6 12.903
35 80 256 8 8 11.103
36 80 320 5 1 4.444
37 80 320 5 2 5.757
38 80 320 5 4 7.761
39 80 320 5 6 12.376
40 80 320 5 8 11.503
41 80 320 8 1 4.550
42 80 320 8 2 6.074
43 80 320 8 4 11.785
44 80 320 8 6 13.218
45 80 320 8 8 11.293
Table 8: Impact of policy on Mixtral 8x22B in Env #2, HumanEval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 32 128 4 4 7.112
2 32 128 4 6 7.921
3 32 128 4 8 7.564
4 32 128 6 4 8.617
5 32 128 6 6 7.901
6 32 128 6 8 7.155
7 32 128 8 4 6.599
8 32 128 8 6 7.913
9 32 128 8 8 7.677
10 32 192 4 4 7.291
11 32 192 4 6 7.083
12 32 192 4 8 4.874
13 32 192 6 4 7.753
14 32 192 6 4 7.733
15 32 192 6 6 7.578
16 32 192 6 8 4.510
17 32 192 8 4 8.536
18 32 192 8 6 6.574
Table 9: Impact of policy on Mixtral 8x22B in Env #2, C-Eval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 16 32 6 4 3.430
2 16 32 6 6 4.510
3 16 32 6 8 4.321
4 16 32 8 4 3.607
5 16 32 8 6 4.230
6 16 32 8 8 4.742
7 32 32 6 4 3.726
8 32 32 6 6 4.977
9 32 32 6 8 4.513
10 32 32 8 4 3.969
11 32 32 8 6 4.233
12 32 32 8 8 3.894
13 32 32 6 4 3.543
14 32 32 6 6 4.647
15 32 32 6 8 4.063
16 32 32 8 4 4.030
17 32 32 8 6 4.231
18 32 32 8 8 3.609
19 16 64 6 4 4.160
20 16 64 6 6 4.510
21 16 64 6 8 3.915
22 16 64 8 4 3.588
Table 10: Impact of policy on Mixtral 8x22B in Env #2, SummEval dataset.
No. Prefill Batch Size Decoding Batch Size Draft batch size Draft max new token Throughput (token/s)
1 16 64 6 4 4.579
2 16 32 6 4 3.711
3 16 32 6 6 3.486
4 16 32 6 8 4.225
5 16 32 8 4 3.862
6 16 32 8 6 3.998
7 16 32 8 8 3.975
8 16 64 6 4 4.529
9 16 64 6 6 5.141
10 16 64 6 8 4.977
11 16 64 8 4 4.546
12 16 64 8 6 4.590
13 16 64 8 8 5.911
A.3.3 GPU Memory Usage
Figure 12: GPU Memory Allocation Overview.
Figure 13: Runtime GPU Memory Monitoring.
We used NVIDIA Nsight [41] to monitor runtime GPU memory usage on the SummEval dataset. As shown in Figure 12, only the parameters essential for target model offloading are retained in memory, while the remaining space is occupied by the draft model and its cache. This aligns with our design rationale: during offloading, it is more efficient to allocate GPU memory to the draft model rather than storing the target model parameters.
Runtime GPU memory monitoring reveals a periodic pattern in the draft model’s memory usage. As shown in the left panel of Figure 13, each cycle lasts approximately 28 seconds, characterized by a gradual increase in memory usage followed by a sharp drop and a 2-second idle window. This aligns with the behavior observed in Figure 7, where the draft model performs computation for 26 seconds and remains idle for 2 seconds awaiting the next batch.
A.3.4 Ablation Study
In addition to the main results, we conducted ablation studies on other datasets. The results are as follows:
Table 11: Ablation study of proposed techniques on HumanEval dataset. The gray tuple denotes a policy (Prefill batch size, decoding batch size, draft batch size, draft max new tokens).
All optimizations No policy search Serial SD No SD
8x7B 34.665 (80, 256, 10, 6) 15.869 (80, 160, 6, 1) 15.005 (80, 256, 32, 6) 16.468 (80, 192, x, x)
8x22B 8.617 (32, 128, 6, 4) 4.510 (32, 192, 6, 8) 5.264 (32, 128, 32, 8) 4.108 (16, 64, x, x)
Table 12: Ablation study of proposed techniques on C-Eval dataset. The gray tuple denotes a policy (Prefill batch size, decoding batch size, draft batch size, draft max new tokens).
All optimizations No policy search Serial SD No SD
8x7B 31.968 (96, 300, 8, 6) 26.126 (96, 288, 8, 2) 21.989 (96, 288, 24, 6) 15.106 (96, 288, x, x)
8x22B 4.977 (32, 32, 6, 6) 3.588 (16, 64, 8, 4) 3.820 (32, 64, 16, 6) 1.812 (32, 64, x, x)
Table 13: Ablation study of proposed techniques on SAMSum dataset. The gray tuple denotes a policy (Prefill batch size, decoding batch size, draft batch size, draft max new tokens).
All optimizations No policy search Serial SD No SD
8x7B 21.109 (100, 300, 6, 4) 12.694 (80, 256, 8, 2) 13.64 (100, 300, 24, 4) 13.072 (80, 256, x, x)
8x22B 4.139 (16, 64, 8, 6) 3.059 (16, 64, 6, 4) 3.544 (16, 64, 16, 6) 2.378(16, 80, x, x)
A.4 Limitation
The main limitation of this paper lies in the fact that speculative decoding is not a consistently reliable method for acceleration. In extreme cases, none of the draft tokens in multiple batches may be accepted, which greatly limits the acceleration effect of SpecOffload.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims are supported by the experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Appendix A.4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All assumptions and complete proofs related to the Planner component are provided in the Appendix A.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our results are reproducible using the attached code. The necessary setup details are given in Appendix A.2.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is available at www.baidu.com. All datasets and models used are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper, together with the Appendix A.2 and instructions in the code, contain information on all substantial details necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The runs for the experiments in the paper have low variance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, we provide sufficient information on the computer resources needed to reproduce the experiments in § 5.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we confirm adherence to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work is focused on efficient LLM Offloading. We do not foresee any particular societal impacts from this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not alter the capabilities of the available models or datasets, but rather provides a more efficient approach to use them on hardware with limited capabilities. Thus, we believe that our paper has a neutral risk impact in this area
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models used are open-source and publicly available.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide a link to the anonymized code, which has documentation provided that explains how to run the experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We did not conduct any crowdsourcing experiments or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We did not conduct any research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is used only for writing, editing, or formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Due to limited GPU memory, existing systems offload model weights to CPU memory, incurring substantial I/O overhead between the CPU and GPU.
-This leads to two major inefficiencies: (1) GPU cores are underutilized, often remaining idle while waiting for data to be loaded; and (2) GPU memory has low impact on performance, as reducing its capacity has minimal effect on overall throughput.
+This leads to two major inefficiencies: (1) GPU cores are underutilized, often remaining idle while waiting for data to be loaded; and (2) GPU memory has a low impact on performance, as reducing its capacity has minimal effect on overall throughput.
 In this paper, we propose SpecOffload, a high-throughput inference engine that embeds speculative decoding into offloading.
@@
 To support this, we carefully orchestrate the interleaved execution of target and draft models in speculative decoding within the offloading pipeline, and propose a planner to manage tensor placement and select optimal parameters.
-Compared to the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×.
+Compared with the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×.
 Our code is available at https://github.com/MobiSense/SpecOffload.
@@
 As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation datageneration , form processing formprocessing , and data wrangling datawrangling .
-These tasks are characterised by LLMs conducting offline inference in batches over large number of tokens.
+These tasks are characterised by LLMs conducting offline inference in batches over a large number of tokens.
 For instance, corporations need to process all archives of financial documentation, whilst individuals want to construct knowledge repositories from accumulated materials.
-In such workloads, higher inference throughput (the number of tokens generated divided by total generation time, token/s) translates into lower total completion time, hence is the key metric.
-Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory constrains.
+In such workloads, higher inference throughput (the number of tokens generated divided by total generation time, token/s) translates into lower total completion time, hence it is the key metric.
+Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory is a major constraint.
 For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) Mixtral8x22b requires at least four NVIDIA H100 (80GB) GPUs.
@@
 There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification quantization ; pruning ; sparsification , etc., which are orthogonal and can be applied on top of offloading.
-Our focus is designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
+Our focus is on designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
 However, the existing offloading work does not utilize GPU resources well.
@@
 As shown in Figure 2, during the decoding phase, the average GPU core utilization of existing methods is only 13% at most.
-This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameters loading.
-To alleviate this issue, existing methods increase the batch size to amortize I/O overhead—by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski .
-However, due to the limitation of GPU memory capacity or CPU computational capabilities, scalability of the batch size remains inherently limited.
+This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameter loading.
+To alleviate this issue, existing methods increase the batch size to amortize I/O overhead by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski .
+However, due to the limitation of GPU memory capacity or CPU computational capabilities, the scalability of the batch size remains inherently limited.
 The maximum batch size achieved by the SOTA scheme in Figure 2 is only 64, insufficient to bridge the huge gap between I/O and GPU latency (even with this batch size, the gap remains over 10×).
@@
 As illustrated in Figure 2, reducing memory usage by over 5.42× for Mixtral 8×7B models led to only a 13% drop in throughput; similarly, a 2.89× reduction for Mixtral 8×22B models caused just a 5% decline.
-This phenomenon arises because the total volume of data to be loaded remains nearly constant, as the model size vastly exceeds the available GPU memory—leaving little room for variation even when accounting for partical parameters that can reside permanently in memory.
+This phenomenon arises because the total volume of data to be loaded remains nearly constant, as the model size vastly exceeds the available GPU memory, leaving little room for variation even when accounting for partical parameters that can reside permanently in memory.
 For instance, generating a single token necessitates loading all (56) FFN layers for the leftmost blue point in Figure 2, compared to 53 layers for the rightmost blue point.
 Consequently, memory reduction precipitates only negligible changes in total inference time and throughput.
-To harnesses GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD).
+To harness GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD).
 SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass blockwise ; FastInference ; ACloseLook .
 SpecOffload embeds SD into the offloading workflow with nearly zero overhead. The key idea lies in the following two aspects:
-  • Computing draft model during GPU core idleness: SD requires draft model to generate multiple candidate tokens in advance.
-Given the substantial idle periods prevalent in existing offloading frameworks, these intervals can be utilized for completing the draft model’s computational tasks.
-  • Storing draft model use "low-yield" GPU memroy: SD requires loading a draft model into memory for draft generation.
+  • Computing draft model during GPU core idleness: SD requires the draft model to generate multiple candidate tokens in advance.
+Given the substantial idle periods prevalent in existing frameworks, these intervals can be utilized for completing the draft model’s computational tasks.
+  • Storing draft model uses "low-yield" GPU memory: SD requires loading a draft model into memory for draft generation.
 We can repurpose "low-yield" memory allocations from existing frameworks to store draft model parameters and its caches instead.
@@
 Our contributions are as follows:
-  • We conduct a quantitative analysis of GPU resource utilization in representative scenarios and identify key limitations in SOTA frameworks——underutilization of GPU cores and marginal utility of GPU memory, thus reveal a novel perspective for enhancing offloading performance.
-  • By designing sub-layer model decomposition and fine-grained scheduling compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
+  • We conduct a quantitative analysis of GPU resource utilization in representative scenarios and identify key limitations in SOTA frameworks—underutilization of GPU cores and marginal utility of GPU memory, thus reveal a novel perspective for enhancing offloading performance.
+  • By designing a sub-layer model decomposition and fine-grained scheduling of compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
   • To evaluate SpecOffload, we benchmark it against HuggingFace Accelerate accelerate , DeepSpeed-FastGen deepspeed , FlexGen Flexgen , and Fiddler fiddler .
@@
 Large language models consist of stacked Transformer layers attentionisallyouneed .
-During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated token and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
+During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated tokens and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
 2.2 Speculative Decoding
@@
 However, traditional single-batch SD is not well-suited for integration with offloading, as the computations of the draft and target models must be executed sequentially.
-As a result, the GPU resources still remain underutilized during the target model’s verification phase.
-By introducing a dual-batch rotation strategy, SpecOffload enable the verifying and drafting to run concurrently, allowing SD to be seamlessly embedded into the offloading pipeline, while better utilizing GPU resources.
+As a result, the GPU resources remain underutilized during the target model’s verification phase.
+By introducing a dual-batch rotation strategy, SpecOffload enables the verifying and drafting to run concurrently, allowing SD to be seamlessly embedded into the offloading pipeline, while better utilizing GPU resources.
 2.3 Offloading in LLM Inference
@@
 As shown in Figure 3, it employs a two-phase architecture where offline tensor placement and online scheduling collectively determine a unified pipeline across GPU and CPU.
-During the offline phase, the target and draft models are deployed across heterogeneous memory hierarchy.
+During the offline phase, the target and draft models are deployed across a heterogeneous memory hierarchy.
 SpecOffload automatically evaluates the hardware ecosystem, measuring CPU and GPU memory capacities, computational performance of CPU and GPU cores, and the bandwidth of data transfer channels between them.
@@
 Motivated by the differing computational demands of the prefill and decoding phases in LLM inference, we introduce the Interleaved Batch Pipeline—a phase-specific pipeline design.
-During the prefill stage, the target model computation dominates the GPU runtime and we perform additional memory management before completion.
+During the prefill stage, the target model computation dominates the GPU runtime, and we perform additional memory management before completion.
 In the decoding stage, we embed speculative decoding into the pipeline by finely interleaving the computations of the two models to fully utilize the GPU cores.
@@
 While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen Flexgen , we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding.
-In order to minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
+To minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
 4.1.2 Decoding Phase
@@
 GPU, IO, and CPU are orchestrated to perform distinct, parallelized tasks.
-From model-level, the decoding phase involves two batches being processed alternately by the target and draft model.
+At model-level, the decoding phase involves two batches being processed alternately by the target and draft model.
 Conventional speculative decoding adopts a single-batch Draft-then-Verify paradigm, as the computations of the draft and target models must be executed sequentially.
-Consequently, the GPU resources left idle during the target model’s verification stage still remain unused, while the system suffers from additional overhead caused by switching between models.
+Consequently, the GPU resources left idle during the target model’s verification stage remain unused, while the system suffers from additional overhead caused by switching between models.
 To overcome these limitations, we propose a dual-batch interleaved design that enables true model-level parallelism.
@@
 This alternating batch rotation continues until the generation is complete.
-This batch interleaving primarily enables parallel execution of drafting and verification.
-From computation-level, the decoding pipeline involves coordination among three threads: GPU computation, CPU computation, and I/O.
+This batch interleaving primarily enables the parallel execution of drafting and verification.
+At computation-level, the decoding pipeline involves coordination among three threads: GPU computation, CPU computation, and I/O.
 The right side of Figure 4 provides an illustrative example.
@@
 Finally, once both the parameters and activations are available on the GPU, the remaining computations are quickly completed.
-After completing all layers of the target model (the end of a round of parallel drafting and verification at model level), the target and draft models exchange their current batches.
+After completing all layers of the target model (the end of a round of parallel drafting and verification at model-level), the target and draft models exchange their current batches.
 Details such as where model parameters are computed, how I/O is handled, and batch configurations are determined in the § 4.2 and § 4.3.
@@
 Although not evaluated, SpecOffload can support other models compatible with Transformers’ model classes.
-Hardware. We evaluate SpecOffload in two different evironments, as shown in Table 2.
+Hardware. We evaluate SpecOffload in two different evironments, as shown in Table 2. Env #2 refers to a cloud-based server.
 Datasets.
@@
 We show the runtime breakdown of Mixtral 8x7B in Env #1 and 8x22B in Env #2 in Table 3 on SummEval dataset.
-We disable overlapping and profile the time used for GPU, I/O and CPU.
+We disable overlapping and profile the time used for GPU, I/O, and CPU.
 Results show that our method effectively overlaps compute and I/O.
@@
 The Draft process writes generated draft tokens to the output shared memory for consumption by the Main process, and reads verified tokens from the input shared memory region.
-When a thread in the Main process reaches the verify stage and requires draft tokens, it waits for the draft model to signal availability, then read from shared memory.
+When a thread in the Main process reaches the verify stage and requires draft tokens, it waits for the draft model to signal availability, then reads from shared memory.
 Upon finishing computation, the draft model signals completion and readiness for the next task using event flags, ensuring proper synchronization and preventing data races or overlap between batches.
@@
 In extreme cases, none of the draft tokens in multiple batches may be accepted, which greatly limits the acceleration effect of SpecOffload.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: All claims are supported by the experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss limitations in Appendix A.4.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: All assumptions and complete proofs related to the Planner component are provided in the Appendix A.1.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Our results are reproducible using the attached code. The necessary setup details are given in Appendix A.2.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Our code is available at www.baidu.com. All datasets and models used are publicly available.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The paper, together with the Appendix A.2 and instructions in the code, contain information on all substantial details necessary to understand the results.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: The runs for the experiments in the paper have low variance.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Yes, we provide sufficient information on the computer resources needed to reproduce the experiments in § 5.1.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Yes, we confirm adherence to the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: Our work is focused on efficient LLM Offloading. We do not foresee any particular societal impacts from this work.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54.
-Justification: This paper does not alter the capabilities of the available models or datasets, but rather provides a more efficient approach to use them on hardware with limited capabilities.
-Thus, we believe that our paper has a neutral risk impact in this area
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All datasets and models used are open-source and publicly available.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We provide a link to the anonymized code, which has documentation provided that explains how to run the experiments.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We did not conduct any crowdsourcing experiments or research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We did not conduct any research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: LLM is used only for writing, editing, or formatting purposes.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 Compared with the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×.
-Our code is available at https://github.com/MobiSense/SpecOffload.
+Our code is available at https://github.com/MobiSense/SpecOffload-public.
 1 Introduction
-As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation datageneration , form processing formprocessing , and data wrangling datawrangling .
+As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation [1], form processing [2], and data wrangling [3].
 These tasks are characterised by LLMs conducting offline inference in batches over a large number of tokens.
@@
 Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory is a major constraint.
-For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) Mixtral8x22b requires at least four NVIDIA H100 (80GB) GPUs.
+For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) [4] requires at least four NVIDIA H100 (80GB) GPUs.
 Offloading is one of the mainstream solutions to memory-constrained inference, transferring most model parameters to more economical, capacious CPU memory and reloading them to GPU memory only when computation demands.
-There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification quantization ; pruning ; sparsification , etc., which are orthogonal and can be applied on top of offloading.
+There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification [5, 6, 7], etc., which are orthogonal and can be applied on top of offloading.
 Our focus is on designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
 However, the existing offloading work does not utilize GPU resources well.
-During offloading, generating each token requires reloading most model parameters from CPU memory to GPU memory for execution deepspeed ; fastinference_offloading ; Klotski .
+During offloading, generating each token requires reloading most model parameters from CPU memory to GPU memory for execution [8, 9, 10].
 Yet I/O speeds substantially lag behind GPU computational capabilities.
@@
 This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameter loading.
-To alleviate this issue, existing methods increase the batch size to amortize I/O overhead by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski .
+To alleviate this issue, existing methods increase the batch size to amortize I/O overhead by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput [11, 10].
 However, due to the limitation of GPU memory capacity or CPU computational capabilities, the scalability of the batch size remains inherently limited.
@@
 When models far exceed GPU memory capacity, throughput remains almost unchanged even when memory usage is greatly reduced during decoding phase.
-We conducted experiments with FlexGen Flexgen , the most effective solution among SOTA approaches.
+We conducted experiments with FlexGen [11], the most effective solution among SOTA approaches.
 FlexGen offloads attention computations to CPU while computing FFN layers on GPU during the decoding phase.
@@
 To harness GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD).
-SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass blockwise ; FastInference ; ACloseLook .
+SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass [12, 13, 14].
 SpecOffload embeds SD into the offloading workflow with nearly zero overhead. The key idea lies in the following two aspects:
@@
 We can repurpose "low-yield" memory allocations from existing frameworks to store draft model parameters and its caches instead.
-For instance, as shown in Figure 2, extracting 17GB of "low-yield" memory allocation provides sufficient capacity for a draft model such as Mistral 7B mistral7b to operate normally within the GPU at a small batch.
+For instance, as shown in Figure 2, extracting 17GB of "low-yield" memory allocation provides sufficient capacity for a draft model such as Mistral 7B [15] to operate normally within the GPU at a small batch.
 To support this, SpecOffload designs a comprehensive framework to better utilize both the computational and memory resources of the GPU.
@@
   • By designing a sub-layer model decomposition and fine-grained scheduling of compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
-  • To evaluate SpecOffload, we benchmark it against HuggingFace Accelerate accelerate , DeepSpeed-FastGen deepspeed , FlexGen Flexgen , and Fiddler fiddler .
+  • To evaluate SpecOffload, we benchmark it against HuggingFace Accelerate [16], DeepSpeed-FastGen [8], FlexGen [11], and Fiddler [17].
 Results demonstrate that our system achieves throughput improvements of 4.69×, 4.71×, 2.54×, and 4.04× over these baselines, respectively.
@@
 2.1 Generative LLM Inference
-Large language models consist of stacked Transformer layers attentionisallyouneed .
-During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated tokens and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
+Large language models consist of stacked Transformer layers [18].
+During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated tokens and the cached KV states [19, 20, 21].
 2.2 Speculative Decoding
@@
 It adheres to a “Draft-then-Verify” framework: at each decoding step, a lightweight draft model initially proposes multiple candidate tokens (e.g., $(\hat{w}_{1},\hat{w}_{2},\hat{w}_{3},\hat{w}_{4}$)), which are collectively verified by a larger target model in a single forward pass.
-Only the valid subset $({w}_{1},{w}_{2})$ is accepted, after which the target model resumes decoding by independently generating the subsequent token $w_{3}$ blockwise ; FastInference ; ACloseLook .
+Only the valid subset $({w}_{1},{w}_{2})$ is accepted, after which the target model resumes decoding by independently generating the subsequent token $w_{3}$ [12, 13, 14].
 This approach enables the target model to generate multiple tokens per inference step.
-To further enhance the efficiency of SD, prior research has predominantly explored two avenues: the design of more effective draft models distillspec ; draft-verify ; onlineSD and draft structures medusa ; specinfer ; specexec .
+To further enhance the efficiency of SD, prior research has predominantly explored two avenues: the design of more effective draft models [22, 23, 24] and draft structures [25, 26, 27].
 However, traditional single-batch SD is not well-suited for integration with offloading, as the computations of the draft and target models must be executed sequentially.
@@
 Offloading is one of the predominant solutions for enabling LLM inference under GPU memory constraints.
-It entails relocating certain model parameters from the expensive, limited GPU memory to the more cost-effective and abundant CPU memory deepspeed ; fiddler ; powerinfer ; moe-infinity ; fastdecode ; fastinference_offloading .
+It entails relocating certain model parameters from the expensive, limited GPU memory to the more cost-effective and abundant CPU memory [8, 17, 28, 29, 30, 9].
 I/O constraints create the primary bottleneck in offloading, as data transfer latency between CPU memory and GPU memory far exceeds the GPU’s computation time.
@@
 In throughput-oriented scenarios, existing methods typically increase the batch size to amortize I/O costs.
-A prevalent strategy involves altering the model’s execution pattern from row-wise to column-wise, allowing each layer’s parameters to be loaded once and reused across multiple batches, thereby reducing the per-layer I/O burden Flexgen ; Klotski .
+A prevalent strategy involves altering the model’s execution pattern from row-wise to column-wise, allowing each layer’s parameters to be loaded once and reused across multiple batches, thereby reducing the per-layer I/O burden [11, 10].
 However, this approach is constrained by the available GPU memory and the overhead I/O of KV cache.
-More recent studies demonstrate that offloading attention computation to the CPU can eliminate KV cache I/O in decoding phase moe-lightning ; neo ; hetegen ; PACT'24 , but in doing so, CPU compute limitations cap batch scalability.
+More recent studies demonstrate that offloading attention computation to the CPU can eliminate KV cache I/O in decoding phase [31, 32, 33, 34], but in doing so, CPU compute limitations cap batch scalability.
 Thus, GPU resources remain significantly underutilized.
@@
 4.1.1 Prefill Phase
-While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen Flexgen , we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding.
+While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen [11], we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding.
 To minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
@@
 Table 2: Dataset Configurations.
-HumanEval humaneval C-Eval ceval SummEval summeval SAMSum samsum
+HumanEval [35] C-Eval [36] SummEval [37] SAMSum [38]
 $S_{avg}$ 157.54 165.46 503.02 168.10
@@
 Implementation.
-We implement SpecOffload on top of Hugging Face Transformers v4.47.1transformers .
+We implement SpecOffload on top of Hugging Face Transformers v4.47.1[39].
 We implement pipeline using multiprocessing with shared memory for inter-process vector communication.
@@
 Models.
-We evaluate SpecOffload using two popular and open-source models: Mixtral-8x7B Mixtral8x7b and Mixtral-8x22B Mixtral8x22b .
+We evaluate SpecOffload using two popular and open-source models: Mixtral-8x7B [40] and Mixtral-8x22B [4].
 They have 46.7B and 141B parameters in bfloat16 precision.
-The draft model for speculative decoding is Mistral-7B mistral7b .
+The draft model for speculative decoding is Mistral-7B [15].
 Although not evaluated, SpecOffload can support other models compatible with Transformers’ model classes.
@@
 We use most common LLM benchmarks with different prompt length distributions and tasks, as shown in Table 2.
-HumanEval dataset humaneval released by OpenAI includes 164 programming problems; C-Eval dataset ceval is a comprehensive Chinese evluation suite includes 13948 questions; SummEval dataset summeval inlcudes 100 news article from the CNN/DailyMail; SAMSum dataset samsum contains about 16k messenger-like conversations with summaries.
+HumanEval dataset [35] released by OpenAI includes 164 programming problems; C-Eval dataset [36] is a comprehensive Chinese evluation suite includes 13948 questions; SummEval dataset [37] inlcudes 100 news article from the CNN/DailyMail; SAMSum dataset [38] contains about 16k messenger-like conversations with summaries.
 Baselines. We compare against 4 baseline systems, all designed to address GPU memory limitations.
-  • Hugging Face Accelerate accelerate supports offloading weights of some layers based on the device map. We use its version 1.5.2. Hereinafter referred to as Accelerate.
-  • DeepSpeed Zero-Inference deepspeed supports offloading the whole weights to the CPU or disk. We use its version 0.16.1. Hereinafter referred to as DeepSpeed.
-  • FlexGen Flexgen employs a "zig-zag" inference schedule to increase throughput.
-  • Fiddler fiddler strategically utilizes both CPU and GPU resources for MoE model inference.
+  • Hugging Face Accelerate [16] supports offloading weights of some layers based on the device map. We use its version 1.5.2. Hereinafter referred to as Accelerate.
+  • DeepSpeed Zero-Inference [8] supports offloading the whole weights to the CPU or disk. We use its version 0.16.1. Hereinafter referred to as DeepSpeed.
+  • FlexGen [11] employs a "zig-zag" inference schedule to increase throughput.
+  • Fiddler [17] strategically utilizes both CPU and GPU resources for MoE model inference.
 Additionally, except for FlexGen, all other baselines natively support Mixtral. We adapted FlexGen to Mixtral while adhering to its original offloading strategy.
@@
 Figure 7: Decoding phase memory consumption.
-We employ NVIDIA Nsight nsight to monitor GPU core utilization and memory consumption during the decoding phase of Mixtral 8×7B in Env #1 on the SummEval dataset.
+We employ NVIDIA Nsight [41] to monitor GPU core utilization and memory consumption during the decoding phase of Mixtral 8×7B in Env #1 on the SummEval dataset.
 As depicted in Figure 7, the average core utilization reaches 58.67%, attains 8.14×, 7.15×, 4.49×, and 8.24× higher than Accelerate, DeepSpeed, FlexGen, and Fiddler, respectively.
@@
 In extreme cases, none of the draft tokens in multiple batches may be accepted, which greatly limits the acceleration effect of SpecOffload.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: All claims are supported by the experimental results.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss limitations in Appendix A.4.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: All assumptions and complete proofs related to the Planner component are provided in the Appendix A.1.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Our results are reproducible using the attached code. The necessary setup details are given in Appendix A.2.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Our code is available at www.baidu.com. All datasets and models used are publicly available.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The paper, together with the Appendix A.2 and instructions in the code, contain information on all substantial details necessary to understand the results.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: The runs for the experiments in the paper have low variance.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Yes, we provide sufficient information on the computer resources needed to reproduce the experiments in § 5.1.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Yes, we confirm adherence to the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: Our work is focused on efficient LLM Offloading. We do not foresee any particular societal impacts from this work.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54.
+Justification: This paper does not alter the capabilities of the available models or datasets, but rather provides a more efficient approach to use them on hardware with limited capabilities.
+Thus, we believe that our paper has a neutral risk impact in this area
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All datasets and models used are open-source and publicly available.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provide a link to the anonymized code, which has documentation provided that explains how to run the experiments.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: We did not conduct any crowdsourcing experiments or research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: We did not conduct any research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: LLM is used only for writing, editing, or formatting purposes.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.